# Measuring AI “Slop” in Text

### v1

@@ 6.4 Training Span Extraction Models
We trained a Qwen-7B reasoning model (DeepSeek-R1-Distill-Qwen-7B; DeepSeek-AI 2025) for slop span extraction. To provide rationales, we first generated silver annotations by prompting GPT-5 with the annotated span and label, asking for concise explanations of the label. We also augment our News and MS Marco data with data from LAMP Chakrabarty et al. (2025a; b), mapping their categories (such as cliche, redundant/exposition) into our slop taxonomy to create a consistent label space. We filter the LAMP dataset to remove the creative writing subsets. We provide details of prompts used, label mappings, and training details in Appendix H.
@@ 7 Discussion
LLMs are often deployed as cheap alternatives for human preference judgments in alignment and evaluation (Bharadwaj et al., 2025), however our findings highlight important limitations. Unlike reasoning tasks where rewards are verifiable, for subjective tasks there is a significant risk of miscalibration. Prior work has documented these issues: Chakrabarty et al. (2025a) and Gooding et al. (2025), for example, show that LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. This leads to low quality text that is often referred to as “slop”.
A recent study from OpenAI Chatterji et al. (2025) shows that almost 50% of ChatGPT usage focuses on writing (28.1%) and information seeking (21.3 %) tasks. To ensure better alignment in such areas, we present the first systematic attempt at qualitatively characterizing “slop” in LLM-generated text. Our findings suggest that Information Quality, Information Utility, and Style Quality are important axes of text evaluation. Further, granular codes within each axis can vary in strength based on the domain, or the purpose of the text. We show that our taxonomy provides a useful framework for assessing writing across domains, beyond accuracy- or reference-based metrics. While overall “slop” judgments are somewhat subjective, our analysis shows that an increase in issues along these axes increases the likelihood of a text being judged as “slop”. We also show that current evaluation practices are not sufficient for automatically measuring “slop”. Existing automatic text metrics fail to capture whether generated text is genuinely useful or well-written relative to “slop”. Neither LLMs-as-judges nor linear models trained over these features are able to fully approximate human assessments of “slop,” however we hope the taxonomy can guide future improvements of LLM-based reward models.
@@ Appendix A Definition Collection
We provide the full survey sent to experts in Table 8. Here, we show the full set of questions and answer options in the survey. Experts explicitly provided permission at the end to share anonymized and aggregate responses to the survey.
@@ Appendix B Additional Survey Findings
Experts provided their years of experience in their reported fields (Table 3(b)), which we report in aggregate in Table 3(a). Most experts had $\geq$3 years of experience, and many had professional experience in the field of NLP or writing. Additionally, experts identified characteristics of “slop” that can appear in human-written text, but all point to qualities that serve a different purpose than those identified for AI-generated text.
@@ Appendix D Consent Forms
Appendix E Annotation Details
@@
A key indicator of “slop” is the relatively low density of information within it. Such texts are often verbose without conveying much information, or contain many generic statements that are broadly applicable. We measure information density in two ways. In the first, we adopt an information-theoretic approach, following Meister et al. (2021). The uniform density hypothesis posits that speakers generally tend towards spreading information uniformly across utterances. We measure token entropy using GPT-2 (Radford et al., 2019), and then evaluate the mean and coefficient of variation. A higher mean indicates an overall lower information density in text, and a higher coefficient of variation indicates less uniformity, both of which can be indicative of “slop”.
@@
Measuring relevance (to a context), similar to factuality, is an active research area. Most methods assume access to a high-quality set of source documents and queries. Relevance, where context is provided, is measured relative to the query and task at hand. In the absence of additional context (e.g., task or domain), relevance can be evaluated on the internal consistency of the passage. “Slop” can comprise content that fails to address the query or task, sometimes subtly. Recently, Clarke & Dietz (2024) showed that GPT-4o cannot reliably act as a replacement for human assessments of relevance for conditional generation. Therefore, we rely on human assessments of relevance with additional context provided where possible.
@@
In non-fiction texts, high-quality text is accurate. LLM “slop”, however, is defined as having “subtle inaccuracies”, introducing hallucinations (“non-existing entities”), or containing fallacious claims. Automatically measuring factuality is an open research problem (e.g., Ramprasad & Wallace 2024; Laskar et al. 2023), and can depend on whether reference (source) documents are available. We rely on human annotations to detect inaccuracies in LLM-written texts in all cases.
@@
Bias in text can refer to a range of topics that might influence the subjectivity of writing, and can span social (Blodgett et al., 2020) or cognitive (Atreides & Kelley, 2024) facets. Unless explicitly prompted to produce an opinion, much of the content in “slop” lacks subjectivity in presenting information (factual or otherwise). Of the expert definitions that mention bias, there is a notable focus on the lack of subjectivity in “slop.” There is often a missing rhetorical point of view when it is otherwise needed, or a lack of appropriate, engaged perspective. For instance, an LLM-generated movie review that simply states facts such as “[…] movie received 3.5 stars and had a small budget” does not provide any of the subjective assessments one expects in a critique.
@@
When defined in the context of “slop”, repetition entails excessive use of the same words or phrases and low diversity in vocabulary and expression. Prior work has looked at measuring semantic (Tevet & Berant, 2020; Namuduri et al., 2025) and lexical repetition (Shaib et al., 2024a). We focus specifically on lexical repetition metrics, measuring the compression ratio over words (CR) and over parts-of-speech (CR: PoS) to capture repetitive phrases and words.
@@
Automatically measuring coherence and fluency in text is difficult (Li et al., 2024; Murugadoss et al., 2024), and may require human assessments to validate. Fluency is the correctness of the written language. Coherence refers to the logical flow and connection between ideas presented in a text. State-of-the-art LLMs that have undergone rounds of post-training and alignment rarely produce text that is completely disfluent. [^12] At least in English; Multilingual assessments may show otherwise. “Slop”, however, can exhibit low coherence (such as poor sentence organization, inconsistencies in argument or narrative, or written in a way that demands significant effort to follow), or subtle disfluencies (e.g., strange turns of phrase, technically correct but unnatural language, or word choices misaligned to the context). We rely on expert human annotations to identify instances of disfluency or incoherence in texts.
@@
Word complexity assesses the vocabulary in a passage relative to the context: “Slop” can contain unnecessary jargon, buzzword-laden content, or can exhibit an overuse of “rare” words (Hovy, 2016). Our evaluation of “slop” is in English texts, so we opt to use established measurements of complexity: Gunning-Fog Index, which measures the years of formal education needed to understand text on a first reading (Gunning, 1952), Flesch-Kincaid Grade Level (Kincaid et al., 1975), measuring the (U.S.) school grade level one needs to understand the text, Flesch Reading Ease (Flesch, 1948), measuring textual difficulty on a 100-point scale where higher scores indicate easier-to-read text. We also measure sentence and word length (Dale & Chall, 1948), as these directly correlate with text complexity.
@@
Appendix F Results by Individual
@@
Appendix G Automated Metrics
@@
Appendix H LLM Evaluations
H.1 Prompting
SYSTEM_PROMPT_SPANS = (
@@ H.1 Prompting
 ’Return a JSON object: { "spans": ["...","..."], "reasoning": "..." }’
SYSTEM_PROMPT_LABEL = (
[[... slop codes and examples ...]]
Task: Is this slop (0 = no, 1 = yes)
H.2 Training
@@
H.3 Data Augmentation (LAMP)

### v2

@@ 6.4 Training Span Extraction Models
We trained a Qwen-7B reasoning model (DeepSeek-R1-Distill-Qwen-7B; DeepSeek-AI 2025) for slop span extraction. To provide rationales, we first generated silver annotations by prompting GPT-5 with the annotated span and label, asking for concise explanations of the label. We also augment our News and MS Marco data with data from LAMP Chakrabarty et al. (2025a; b), mapping their categories (such as cliche, redundant/exposition) into our slop taxonomy to create a consistent label space. We filter the LAMP dataset to remove the creative writing subsets. We provide details of prompts used, label mappings, and training details in Appendix L.
@@ 7 Discussion
LLMs are often deployed as cheap alternatives for human preference judgments in alignment and evaluation (Bharadwaj et al., 2025), however our findings highlight important limitations. Unlike reasoning tasks where rewards are verifiable, for subjective tasks there is a significant risk of miscalibration. Prior work has documented these issues: Chakrabarty et al. (2025a) and Gooding et al. (2025), for example, show that LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. This leads to low quality text that is often referred to as “slop".
A recent study from OpenAI (Chatterji et al., 2025b) shows that almost 50% of ChatGPT usage focuses on writing (28.1%) and information seeking (21.3 %) tasks. To ensure better alignment in such areas, we present the first systematic attempt at qualitatively characterizing “slop” in LLM-generated text. Our findings suggest that Information Quality, Information Utility, and Style Quality are important axes of text evaluation. Further, granular codes within each axis can vary in strength based on the domain, or the purpose of the text. We show that our taxonomy provides a useful framework for assessing writing across domains, beyond accuracy- or reference-based metrics. While overall “slop” judgments are somewhat subjective, our analysis shows that an increase in issues along these axes increases the likelihood of a text being judged as “slop”. We also show that current evaluation practices are not sufficient for automatically measuring “slop”. Existing automatic text metrics fail to capture whether generated text is genuinely useful or well-written relative to “slop”. Neither LLMs-as-judges nor linear models trained over these features are able to fully approximate human assessments of “slop,” however we hope the taxonomy can guide future improvements of LLM-based reward models. While our analysis focuses on text written in English, we hope the framework introduced here can support future analyses of other languages.
@@ Appendix A Definition Collection
We provide the full survey sent to annotators in Table 8. Here, we show the full set of questions and answer options in the survey. Annotators explicitly provided permission at the end to share anonymized and aggregate responses to the survey.
@@ Appendix B Additional Survey Findings
Experts provided their years of experience in their reported fields (Table 4(b)), which we report in aggregate in Table 4(a). Most annotators had $\geq$3 years of experience, and many had professional experience in the field of NLP or writing. Additionally, experts identified characteristics of “slop” that can appear in human-written text, but all point to qualities that serve a different purpose than those identified for AI-generated text.
@@ Appendix D Consent Forms
Appendix E Expert Annotator Details
Here, we provide details of the backgrounds and expertise of our selected annotators. Our analyses are limited to English texts so we hired annotators who are fluent in English. All annotators are native English speakers from North America. Each annotator had between 15 and 30 years of experience working as professional copy-editors and writers, in Education, Publication, and Business fields. Annotators have worked additionally as writers for print and online professional publications, and as educators teaching writing.
Appendix F Annotation Details
@@
A key indicator of “slop” is the relatively low density of information within it. Such texts are often verbose without conveying much information, or contain many generic statements that are broadly applicable. We measure information density in two ways. In the first, we adopt an information-theoretic approach, following Meister et al. (2021). The uniform density hypothesis posits that speakers generally tend towards spreading information uniformly across utterances. We measure token entropy using GPT-2 , and then evaluate the mean and coefficient of variation. A higher mean indicates an overall lower information density in text, and a higher coefficient of variation indicates less uniformity, both of which can be indicative of “slop”.
@@
Measuring relevance (to a context), similar to factuality, is an active research area. Most methods assume access to a high-quality set of source documents and queries. Relevance, where context is provided, is measured relative to the query and task at hand. In the absence of additional context (e.g., task or domain), relevance can be evaluated on the internal consistency of the passage. “Slop” can comprise content that fails to address the query or task, sometimes subtly. Recently, Clarke and Dietz (2024) showed that GPT-4o cannot reliably act as a replacement for human assessments of relevance for conditional generation. Therefore, we rely on human assessments of relevance with additional context provided where possible.
@@
In non-fiction texts, high-quality text is accurate. LLM “slop”, however, is defined as having “subtle inaccuracies”, introducing hallucinations (“non-existing entities”), or containing fallacious claims. Automatically measuring factuality is an open research problem (e.g., Ramprasad and Wallace 2025), and can depend on whether reference (source) documents are available. We rely on human annotations to detect inaccuracies in LLM-written texts in all cases.
@@
Bias in text can refer to a range of topics that might influence the subjectivity of writing, and can span social (Blodgett et al., 2020) or cognitive (Atreides and Kelley, 2024) facets. Unless explicitly prompted to produce an opinion, much of the content in “slop” lacks subjectivity in presenting information (factual or otherwise). Of the expert definitions that mention bias, there is a notable focus on the lack of subjectivity in “slop.” There is often a missing rhetorical point of view when it is otherwise needed, or a lack of appropriate, engaged perspective. For instance, an LLM-generated movie review that simply states facts such as “[…] movie received 3.5 stars and had a small budget” does not provide any of the subjective assessments one expects in a critique.
@@
When defined in the context of “slop”, repetition entails excessive use of the same words or phrases and low diversity in vocabulary and expression. Prior work has looked at measuring semantic (Tevet and Berant, 2020; Namuduri et al., 2025) and lexical repetition (Shaib et al., 2024a). We focus specifically on lexical repetition metrics, measuring the compression ratio over words (CR) and over parts-of-speech (CR: PoS) to capture repetitive phrases and words.
@@
Automatically measuring coherence and fluency in text is difficult (Li et al., 2024; Murugadoss et al., 2025), and may require human assessments to validate. Fluency is the correctness of the written language. Coherence refers to the logical flow and connection between ideas presented in a text. State-of-the-art LLMs that have undergone rounds of post-training and alignment rarely produce text that is completely disfluent. [^14] At least in English; Multilingual assessments may show otherwise. “Slop”, however, can exhibit low coherence (such as poor sentence organization, inconsistencies in argument or narrative, or written in a way that demands significant effort to follow), or subtle disfluencies (e.g., strange turns of phrase, technically correct but unnatural language, or word choices misaligned to the context). We rely on expert human annotations to identify instances of disfluency or incoherence in texts.
@@
Word complexity assesses the vocabulary in a passage relative to the context: “Slop” can contain unnecessary jargon, buzzword-laden content, or can exhibit an overuse of “rare” words (Hovy, 2016). Our evaluation of “slop” is in English texts, so we opt to use established measurements of complexity: Gunning-Fog Index, which measures the years of formal education needed to understand text on a first reading (Gunning, 1952), Flesch-Kincaid Grade Level (Kincaid et al., 1975), measuring the (U.S.) school grade level one needs to understand the text, Flesch Reading Ease (Flesch, 1948), measuring textual difficulty on a 100-point scale where higher scores indicate easier-to-read text. We also measure sentence and word length as these directly correlate with text complexity.
@@
Appendix G Human-LLM Span Overlap (Qualitative Assessment
Figure 8 shows the top ranked trigrams, and their categorization along the “slop” themes by colour. o3-mini tends to overly reason about information density relative to human-assigned labels.
Figure 8: Tri-grams extracted from o3-mini rationales over highlighted “slop” spans.
Appendix H Results by Domain and Topic
In the News domain, we further assess the distribution of “slop” labels stratified by the source of the article (e.g., Discover, Wall Street Journal). Figure 9 shows the distribution of labels from annotators across each News source. We find that the categories are roughly represented similarly across sources (e.g., Style Quality codes annotated at a much higher rate relative to Information Quality within each source). We include the source counts in Table 13.
Publication Number
National Geographic 41
Smithsonian Magazine 38
New York Times 37
Wall Street Journal 31
Discover 29
Readers Digest 24
Associated Press 22
Scientific American 19
Reader’s Digest 7
Table 13: Number of articles for each News source.
Figure 9: “Slop” axis label prevalence stratified by news source.
Appendix I Results by Individual
@@
Appendix J Automated Metrics
@@
Appendix K Character-level Precision and Recall
We provide the pseudo-code for calculating character-level precision and recall for span overlap in Algorithm 1. We note that this can be modified to calculate word-level overlap, and empirically find our conclusions hold when using both character- and word-level evaluations.
Algorithm 1 Span-level Precision, Recall, and F1 Computation
@@
[…]
3:for each row in dfdf do
4:  $text\leftarrow row.\text{text}$
5:  $gold\_spans\leftarrow\text{parse\_spans}(row.\text{gold})$
6:  $gold\_mask\leftarrow\text{mark\_characters}(|text|,gold\_spans)$
7:  append gold_maskgold\_mask to gold_masks
8:  $pred\_spans\leftarrow\text{parse\_spans}(row.\text{pred})$
9:  $pred\_mask\leftarrow\text{mark\_characters}(|text|,pred\_spans)$
10:  append pred_maskpred\_mask to pred_masks
11:end for
[…]
20:return all computed metrics
Appendix L LLM Evaluations
L.1 Training
@@
L.2 Data Augmentation (LAMP)
@@
L.3 Prompting
SYSTEM_PROMPT_SPANS = (
[…]
 ’Return a JSON object: { "spans": ["...","..."], "reasoning": "..." }’
SYSTEM_PROMPT_LABEL = (
[[... slop codes and examples ...]]
Task: Is this slop (0 = no, 1 = yes)